# An Exploratory Case Study of Query Plan Representations

### v1

@@ 3.4. Properties
Classification. We identified four categories of properties, as shown in the right part of Table 2. On average, every DBMS defines 30 properties. PostgreSQL has the most properties, since it includes many fine-grained properties. For example, it defines three properties to show the status of parallel computations: worker number, worker launched, and worker planned, while other DBMSs provide at most one property for parallel computation.
Cardinality. The Cardinality category consists of the numeric properties that denote the estimated data size returned by operations. The properties in this category can be associated with operations of any category or the serialized query plan as a whole. In 1, the properties rows and width belong to the Cardinality category and represent the estimated number of returned rows and width. These estimates are derived from statistical information (Ioannidis, 2003) that DBMSs collect, such as the total number of rows and maximum values. Query plans with lower estimated cardinalities are more likely to be selected for execution during cost-based query optimization. Some DBMSs, such as MySQL, provide more fine-grained information about the number of rows that are read and returned. As a lightweight DBMS, SQLite uses simple heuristics to estimate cardinalities, and omits properties in the Cardinality category.
@@
Configuration. The Configuration category consists of the properties that configure the operations’ parameters, and their values are configuration options which are usually strings or boolean values. The properties in the Configuration category can be associated with operations in any category or the serialized query plan, and are typically specific to operations. In 1, PostgreSQL’s properties Group Key, Set Cond, Sort Key, Recheck Cond, Index Cond, Filter are in the Configuration category and are specific to the associated operations to show the keys used to group, the condition to join, the key to sort, the condition to check, the index condition, and the predicate to exclude data, respectively. SQLite’s property USING INDEX denotes the index condition.
Status. The Status category consists of the properties of run-time status, and their values are runtime metrics which are usually strings or numbers. These properties can be associated with operations in any category or the serialized query plans, and typically differ depending on the operations they are attached to. In 1, the property Workers Planned is in the Status category and shows the number of available computing nodes to execute the associated operation. As another example, TiDB defines the property taskType to show the name of the computing nodes that the operation is assigned to execute. The properties in the Status category show running status, and are determined by the execution environment, while the properties in the Parameters are usually decided by queries. The properties in both Status and Configuration categories are customized, and thus are typically different across DBMSs, while the properties in other categories share similar semantics or functionalities across DBMSs.
3.5. Formats
Table 3. The officially supported formats of query plans.
@@ 3.5. Formats
Table 3 shows the different formats of query plan representations. Overall, DBMSs support more formats in the natural category rather than the structured category, suggesting that DBMSs prioritize readability over machine processing. Due to the lack of a standard, none of the formats is supported by all DBMSs. For the same DBMS, the formats in the natural category usually include less information than the formats in the structured category. For example, in 1, the property Parent Relationship represents how the associated operation passes data to another operation. This property is ignored in the text format of the natural category, but is shown in the JSON format of the structured category. We provide more details of each format as follows.
@@
Structured category. The structured category includes the JSON, XML, and YAML formats. These formats are standardized and widely used for exchanging data (International, 2013; Bosak, 1997). JSON is more widely supported by DBMSs than other structured formats, and PostgreSQL supports all structured formats. Structured formats are not supported by some DBMSs, such as InfluxDB, SQLite, and SparkSQL. SQLite can output a structured format of bytecode, which consists of low-level instructions, not operations and properties, so we do not consider it as a structured query plan representation.
Table 4. Third-party visualization tools for query plans.
@@
Postgres Explain Visualizer 2 (Website, 2023g) PostgreSQL Open-source
pgmustard (Website, 2023f) PostgreSQL Commercial
pganalyze (Website, 2023e) PostgreSQL Commercial
ApexSQL (Website, 2023a) SQL Server Commercial
Plan Explorer (Orthey et al., 2020) SQL Server Commercial
Azure Data Studio (Website, 2023b) SQL Server Commercial
Dbvisualizer (Website, 2023c) MySQL, PostgreSQL, SQL Server Commercial
Visualization. Apart from the graph formats in official IDEs, third-party visualization tools show query plans based on structured formats to enhance the readability of query plans. Table 4 shows the visualization tools we found for the studied DBMSs. Six of the seven tools are commercial, suggesting the value of understanding query plan representations for developers. Building these tools requires non-trivial effort, because a tool is specific to a DBMS.
4. Unified Query Plan Representation
Our study in section 3 shows that query plan representations share the same conceptual basis, which is why we propose a unified query plan representation that is:
  1. (1)
@@ 4. Unified Query Plan Representation
  2. (2)
    general, to support various DBMSs we studied,
  3. (3)
@@
Figure 2. Railroad diagrams of the symbols tree and node.
4.1. Design
To define and illustrate the unified query plan representation, we adopted Extend Backus Naur Form (EBNF) (Pattis, 1994), which is a metasyntax notation to express context-free grammars. 2 shows the unified query plan representation in EBNF. Following our study in section 3, the unified representation includes the three identified conceptual components of different categories. We define plan as a tree that can have plan-associated properties. Within the tree, a node is defined as an operation and zero or multiple operation-associated properties. operations and properties are key-value pairs including corresponding categories, identifiers, and values. In Figure 2, we also visualized the EBNF into a railroad diagram for an intuitive explanation. To make it concise, we only show the symbols tree and node, which express the high-level structure.
We use a unified naming convention to denote operations and properties in the unified query plan representation. A unified naming convention increases the readability and consistency of the representation while avoiding name collisions. section 3 shows that various operations and properties share similar semantics, so we mapped DBMS-specific names of operations and properties to unified names. For example, we mapped the operation name Seq Scan in PostgreSQL, Table Scan in SQL Server, and TableFullScan in TiDB to Full Table Scan.
4.2. Analysis
We qualitatively analyze whether this design achieves three goals.
Completeness. The unified query plan representation includes the three conceptual components: operations, properties, and formats, which we identified in section 3. Operations and properties are included in the tree of the unified representation, and the unified representation can be serialized into other standard formats, such as JSON and XML, which are used in query plan representations.
@@ 4.2. Analysis
Extensibility. The definitions of operations, properties, and categories in the unified query plan representation can be extended or shrunk while keeping forward and backward compatibility. Forward compatibility refers to allowing a system accepts input intended for a later version of itself, and backward compatibility refers to allowing a system accepts input intended for an older version of itself (Wallace, 2007). Both evaluate whether the applications based on the unified query plan representation still work if we update the representation to support more or different DBMSs. To keep forward compatibility, we can add more categories by expanding operation_category to include more category names, and add more operations and properties whose names comply with the definition of the keyword at line 11 at 2. An existing application still can parse the revised representation by ignoring the newly added categories, operations, and properties, or handle them in a generic way (e.g., a visualization tool could represent unknown operations using a generic visual shape). For backward compatibility, similarly, an existing application still can parse the old version of the representation, whose categories, operations, and properties are included in the new version of the representation.
5. Applications
We implemented the prototype of a reusable library, UPlan, to maintain the unified query plan representation. While our main motivation is facilitating the implementation of automated testing approaches, the unified representation enables also other applications of visualization and benchmarking:
A.1 Testing.:
@@ 5. Applications
DBMSs. For A.2 and A.3, we used MongoDB, MySQL, Neo4j, PostgreSQL, and TiDB, because they support the JSON format of query plans. JSON is the most widely supported structural format, and it typically includes more detailed information than other formats. For example, JSON may include read and optimization costs, but the text format often only includes overall cost, as it is typically optimized for human readability. Within these DBMSs, we used MySQL, PostgreSQL, and TiDB for A.1 as they are supported by SQLancer. We used the same versions of DBMSs that we studied in the Table 1.
Data set. For A.1, we used SQLancer to generate test cases. For A.2 and A.3, we collected the serialized query plans of the queries from the TPC-H benchmark suite (Website, 2023h) on the five DBMSs. TPC-H comprises 8 tables and 22 queries for relational DBMSs: MySQL, PostgreSQL, and TiDB in our experiments. MongoDB and Neo4j require manual effort to adapt TPC-H as it was not designed for non-SQL languages and non-relational data models. MongoDB is based on a document data model, which lacks support for join operations, so we embedded all entities in one document and rewrote queries 1, 3, and 4 in the MongoDB Query Language (MQL), following a tutorial (Website, 2024c). For Neo4j, which is based on a graph data model, we mapped nodes into rows of the relational data model and edges into foreign keys of the relational data model following another tutorial (Website, 2024d), and rewrote queries 1–14, 16–19 using the Cypher Query Language (CQL) following an example (Website, 2024a).
@@ A.1 Testing
We show how UPlan allows implementing the testing methods QPG (Ba and Rigger, 2023) and CERT (Ba and Rigger, 2024) in a DBMS-agnostic way. QPG is a test case generation approach that is guided by query plans; specifically, it mutates a database if no new query plans have been observed for a specific number of randomly generated queries, aiming to subsequently exercise new query plans, and thus exploring “interesting” behaviors. In terms of implementation, evaluating whether a query plan is structurally different from another requires ignoring unstable information, such as random identifiers and the estimated cost in query plans. CERT is a test oracle for finding performance issues by comparing estimated cardinalities, which have to be extracted from query plans. Both methods were implemented in SQLancer, which is a popular and widely used tool for automatically testing DBMSs. The original implementations used DBMS-specific and error-prone methods, such as string matching and substitution, which had to be implemented for every DBMS that was supported. Both approaches support relational DBMSs, but were not applied to the popular open-source DBMSs MySQL and PostgreSQL, as additional DBMS-specific parsers would have been required. Based on UPlan, we implemented general parsers for both methods to support all UPlan-compatible DBMSs. We did not consider Mozi, because its source code is not public.
@@
To evaluate UPlan on QPG and CERT, we applied the general parser to MySQL, PostgreSQL, and TiDB. We ran our revised versions of QPG and CERT for 24 hours and found 17 unique and previously unknown bugs, as shown in Table 5. Developers confirmed 16 of 17 bugs and fixed two bugs. The one bug in PostgreSQL was still waiting for the response from developers, so we did not report more bug reports to avoid burdening developers. 11 of 17 bugs are Critical, Serious, or Major, which represent that the bugs can significantly affect the systems. Additionally, CERT found hundreds of potential bug-inducing test cases in 24 hours, but it is challenging to distinguish their uniqueness, which requires developers’ expertise. To avoid burdening developers, we will report them after the above 17 bugs have been fixed. 3 shows a bug in MySQL found by the test case generated by QPG with UPlan. Note that we identified this bug by the test oracle Ternary Logic Partitioning (TLP) (Rigger and Su, 2020); however, for presentation, we simplified the bug-inducing test case by demonstrating that the same query returns different results in lines 4 and 6 depending on whether the index exists. The cause of the bug was an incorrect table look-up due to the index inserted by the SQL statement in line 5. Using UPlan, we were able to apply QPG to MySQL easily, which was previously incompatible and untested by QPG. This enabled us to find this bug in MySQL.
@@
Table 5. Previously unknown and unique bugs found by QPG with UPlan.
@@ A.2 Visualization
We implemented a visualization tool for serialized query plans by modifying PEV2 (Website, 2023g), a customized query plan visualization tool for PostgreSQL, to use the unified query plan representation. We modified its parser to support identifying the unified query plan representation, and updated its definitions of visualized elements, such as operation names.
Figure 3. Visualized unified query plans of query 1 from TPC-H benchmark.
Using UPlan, our customized PEV2 supports query plan visualization of the five DBMSs in the same implementation by modifying around 800 lines of code. Suppose implementing a new visualization tool for another DBMS also requires at least an additional 800 lines of code, then it would require $800*(5-1)=3,200$ lines of code to support the five DBMSs based on PEV2 without the unified query plan representation. UPlan reduces the effort of building a new visualization tool for serialized query plans and unifies the query plan visualizations across DBMSs.

### v3

@@ III-D Properties
Classification. We identified four categories of properties, as shown in the right part of Table II. On average, every DBMS defines 30 properties. PostgreSQL has the most properties, since it includes many fine-grained properties. For example, it defines three properties to show the status of parallel computations: worker number, worker launched, and worker planned, while other DBMSs provide at most one property for parallel computation. We give a detailed explanation of each property category as follows.
Cardinality. The Cardinality category consists of the numeric properties that denote the estimated data size returned by operations. The properties in this category can be associated with operations of any category or the serialized query plan as a whole. In 1, the properties rows and width belong to the Cardinality category and represent the estimated number of returned rows and width. These estimates are derived from statistical information [32] that DBMSs collect, such as the total number of rows and maximum values. Query plans with lower estimated cardinalities are more likely to be selected for execution during cost-based query optimization. Some DBMSs, such as MySQL, provide more fine-grained information about the number of rows that are read and returned. As a lightweight DBMS, SQLite uses simple heuristics to estimate cardinalities, and omits properties in the Cardinality category.
@@
Configuration. The Configuration category consists of the properties that configure the operations’ parameters, and their values are configuration options which are usually strings or boolean values. The properties in the Configuration category can be associated with operations in any category or the serialized query plan, and are typically specific to operations. In 1, PostgreSQL’s properties Group Key, Set Cond, Sort Key, Recheck Cond, Index Cond, Filter are in the Configuration category and are specific to the associated operations to show the keys used to group, the condition to join, the key to sort, the condition to check, the index condition, and the predicate to exclude data, respectively. SQLite’s property USING COVERING INDEX denotes the index condition.
Status. The Status category consists of the properties of run-time status, and their values are runtime metrics which are usually strings or numbers. These properties can be associated with operations in any category or the serialized query plans, and typically differ depending on the operations they are attached to. In 1, the property Workers Planned is in the Status category and shows the number of available computing nodes to execute the associated operation. As another example, due to the distributed architecture, TiDB defines the property taskType to show the nodes that the operation is assigned to execute on. The properties in the Status category show running status, and are determined by the execution environment, while the properties in the Parameters are usually decided by queries. The properties in both Status and Configuration categories are customized, and thus are typically different across DBMSs, while the properties in other categories share similar semantics or functionalities across DBMSs.
III-E Formats
TABLE III: The officially supported formats of query plans.
@@ III-E Formats
Table III shows the different formats of query plan representations. Overall, DBMSs support more formats in the natural category rather than the structured category, suggesting that DBMSs prioritize readability over machine processing. Due to the lack of a standard, none of the formats is supported by all DBMSs. For the same DBMS, the formats in the natural category usually include less information than the formats in the structured category. For example, in 1, the property Parent Relationship represents how the associated operation passes data to another operation. This property is ignored in the text format of the natural category, but is shown in the JSON format of the structured category. We provide more details of each format as follows.
@@
Structured category. The structured category includes the JSON, XML, and YAML formats. These formats are standardized and widely used for exchanging data [33, 34]. JSON is more widely supported by DBMSs than other structured formats, and PostgreSQL supports all structured formats. Structured formats are not supported by some DBMSs, such as InfluxDB, SQLite, and SparkSQL. SQLite can output a structured format of bytecode, which consists of low-level instructions, not operations and properties, so we do not consider it as a structured query plan representation.
TABLE IV: Third-party visualization tools for query plans.
@@
Postgres Explain Visualizer 2 [35] PostgreSQL Open-source
pgmustard [36] PostgreSQL Commercial
pganalyze [37] PostgreSQL Commercial
ApexSQL [38] SQL Server Commercial
Plan Explorer [39] SQL Server Commercial
Azure Data Studio [40] SQL Server Commercial
Dbvisualizer [41] MySQL, PostgreSQL, SQL Server Commercial
Visualization. Apart from the graph formats in official IDEs, third-party visualization tools show query plans based on structured formats to enhance the readability of query plans. Table IV shows the visualization tools we found for the studied DBMSs. Six of the seven tools are commercial, suggesting the value of understanding query plan representations for developers. Building these tools requires non-trivial effort, because a tool is specific to a DBMS.
IV Unified Query Plan Representation
Our study in Section III shows that query plan representations share the same conceptual basis, which is why we propose a unified query plan representation that is:
  1. 1.
@@ IV Unified Query Plan Representation
  2. 2.
    general, to support various DBMSs we studied, and,
  3. 3.
@@
IV-A Design
To define and illustrate the unified query plan representation, we adopted Extend Backus Naur Form (EBNF) [42], which is a metasyntax notation to express context-free grammars. 2 shows the unified query plan representation in EBNF. Following our study in Section III, the unified representation includes the three identified conceptual components of different categories. We define plan as a tree that can have plan-associated properties. Within the tree, a node is defined as an operation and zero or multiple operation-associated properties. operations and properties are key-value pairs including corresponding categories, identifiers, and values.
We use a unified naming convention to denote operations and properties in the unified query plan representation. A unified naming convention increases the readability and consistency of the representation while avoiding name collisions. Section III shows that various operations and properties share similar semantics, so we mapped DBMS-specific names of operations and properties to unified names. For example, we mapped the operation name Seq Scan in PostgreSQL, Table Scan in SQL Server, and TableFullScan in TiDB to Full Table Scan.
IV-B Analysis
We analyze whether this design achieves the three goals.
Completeness. The unified query plan representation includes the three conceptual components: operations, properties, and formats, which we identified in Section III. Operations and properties are included in the tree of the unified representation, and the unified representation can be serialized into other standard formats, such as JSON and XML, which are used in query plan representations.
@@ IV-B Analysis
Extensibility. The definitions of operations, properties, and categories in the unified query plan representation can be extended or shrunk while keeping forward and backward compatibility. Forward compatibility refers to allowing a system to accept input intended for a later version of itself, and backward compatibility refers to allowing a system to accept input intended for an older version of itself [43]. Both evaluate whether the applications based on the unified query plan representation still work if we update the representation to support more or different DBMSs. To keep forward compatibility, we can add more categories by expanding operation_category to include more category names, and add more operations and properties whose names comply with the definition of the keyword at line 11 at 2. The required effort is minimal, as it involves only adding keyword definitions for any new operation or property in a specific DBMS, while the rest of the unified query plan representation remains unaffected. An existing application still can parse the revised representation by ignoring the newly added categories, operations, and properties, or handling them in a generic way (e.g., a visualization tool could represent unknown operations using a generic visual shape). For backward compatibility, similarly, an existing application still can parse the old version of the representation, whose categories, operations, and properties are included in the new version of the representation, so no maintenance effort is required.
An Example of Extensibility. We assume that one of the DBMSs, PostgreSQL, introduces a new operation in query plans to implement Large Language Model (LLM)-based joining [44]. The UPlan developers would require minimal effort to accommodate this feature by adding the keyword LLM Join for the new operation, without impacting the rest of the unified query plan representation. Similarly, to deprecate this feature, we would remove the keyword in the grammar. These changes would be both forward and backward-compatible to the applications of UPlan. Suppose a visualization tool supports UPlan to visualize query plans for our studied DBMSs. It is forward-compatible as UPlan provides the information of LLM Join and the visualization tool can visualize it without any modification. It is also backward-compatible as older versions of UPlan would still be supported by the visualization tool, as any older formats would have at most as many operations as the currently-supported format.
V Applications
We implemented the prototype of a reusable library, UPlan, to maintain the unified query plan representation. While our main motivation is facilitating the implementation of automated testing approaches, the unified representation also enables other applications of visualization and benchmarking:
A.1 Testing.
@@ V Applications
DBMSs. For A.2 and A.3, we used MongoDB, MySQL, Neo4j, PostgreSQL, and TiDB, because they support the JSON format of query plans. JSON is the most widely supported structural format, and it typically includes more detailed information than other formats. For example, JSON may include read and optimization costs, but the text format often only includes overall cost, as it is typically optimized for human readability. Within these DBMSs, we used MySQL, PostgreSQL, and TiDB for A.1 as they are supported by SQLancer. We used the same versions of DBMSs that we studied in the Table I.
Data set. For A.1, we used SQLancer to generate test cases. For A.2 and A.3, we used the benchmarks TPC-H [45] for the five DBMSs, YCSB [46] for MongoDB, and WDBench [47] for Neo4j. TPC-H is one of the most established benchmarks consisting of business-oriented ad-hoc queries and concurrent data modification for evaluating relational DBMSs. It comprises 8 tables and 22 queries for relational DBMSs: MySQL, PostgreSQL, and TiDB in our experiments. For MongoDB and Neo4j, TPC-H requires manual effort to execute as it was not designed for non-SQL languages and non-relational data models. MongoDB is based on a document data model, which lacks support for join operations, so we embedded all entities in one document and rewrote queries 1, 3, and 4 in the MongoDB Query Language (MQL), following a tutorial [48]. For Neo4j, which is based on a graph data model, we mapped nodes into rows of the relational data model and edges into foreign keys of the relational data model following another tutorial [49], and rewrote queries 1–14, 16–19 using the Cypher Query Language (CQL) following an example [50]. WDBench evaluates graph DBMSs and consists of 2623 queries. YCSB is used for NoSQL DBMSs and dynamically generates workloads including updates and queries. We used both benchmarks to further explore the query plans in MongoDB and Neo4j.
@@ A.1 Testing
We show how UPlan allows implementing the testing methods QPG [4] and CERT [5] in a DBMS-agnostic way. QPG is a test case generation approach that is guided by query plans; specifically, it mutates a database if no new query plans have been observed for a specific number of randomly generated queries, aiming to subsequently exercise new query plans, and thus exploring “interesting” behaviors. In terms of implementation, evaluating whether a query plan is structurally different from another requires ignoring unstable information, such as random identifiers and the estimated cost in query plans. CERT is a test oracle for finding performance issues by comparing estimated cardinalities, which have to be extracted from query plans. Both methods were implemented in SQLancer, which is a popular and widely used tool for automatically testing DBMSs. The original implementations used DBMS-specific and error-prone methods, such as string matching and substitution, which had to be implemented for every DBMS that was supported. Both approaches support relational DBMSs, but were not applied to the popular open-source DBMSs MySQL and PostgreSQL, as additional DBMS-specific parsers would have been required. Based on UPlan, we implemented general parsers for both methods to support all UPlan-compatible DBMSs. We did not consider Mozi, because its source code is not public.
Figure 2: The architecture of using UPlan for QPG and CERT. Grey font means the implementation without UPlan, and $\checkmark$ refers to the implementation with UPlan. QPG/CERT stops when observing an unexpected result or query plan.
Figure 2 shows the architecture of using UPlan for QPG and CERT. Both QPG and CERT iteratively execute queries in target DBMSs and retrieve query plans for generating the next query and examining cardinalities respectively. In this example, suppose both QPG and CERT execute the query EXPLAIN SELECT * FROM t0 WHERE t0<5 in MySQL, PostgreSQL, and TiDB, the query plans returned by three DBMSs are significantly different. Without UPlan, QPG and CERT require DBMS-specific implementation to parse the raw query plans. With UPlan, we converted raw query plans into the unified query plan, and QPG and CERT require only a single implementation based on UPlan. In this example, TiDB’s plan is converted into two operations, while PostgreSQL’s and MySQL’s plans are converted into Producer->Full Table Scan only as TiDB requires Executor->Collect to receive data from other nodes in a distributed system.
@@
To evaluate UPlan on QPG and CERT, we applied the general parser to MySQL, PostgreSQL, and TiDB. We ran our revised versions of QPG and CERT for 24 hours and found 17 unique and previously unknown bugs, as shown in Table V. Developers confirmed 16 of the 17 bugs and fixed two bugs. The one bug in PostgreSQL was still waiting for the response from developers, so we did not report more bug reports to avoid burdening developers. 11 of 17 bugs are Critical, Serious, or Major, which represent that the bugs can significantly affect the systems. Additionally, CERT found hundreds of potential bug-inducing test cases in 24 hours, but it is challenging to distinguish their uniqueness, which requires developers’ expertise. 3 shows a bug in MySQL found by the test case generated by QPG with UPlan. Note that we identified this bug by the test oracle Ternary Logic Partitioning (TLP) [51]; however, for presentation, we simplified the bug-inducing test case by demonstrating that the same query returns different results in lines 4 and 6 depending on whether the index exists. The cause of the bug was an incorrect table look-up due to the index inserted by the SQL statement in line 5. Using UPlan, we were able to apply QPG to MySQL easily, which was previously incompatible and untested by QPG. This enabled us to find this bug in MySQL.
@@
TABLE V: Previously unknown and unique bugs found by QPG with UPlan.
@@ A.2 Visualization
We implemented a visualization tool for serialized query plans by modifying PEV2 [35], a customized query plan visualization tool for PostgreSQL, to use the unified query plan representation. We modified its parser to support identifying the unified query plan representation, and updated its definitions of visualized elements, such as operation names.
Figure 3: Visualized unified query plans of query 1 from TPC-H benchmark.
UPlan significantly reduces the effort to build visualization tools on query plans. According to the Git repository, developers of PEV2 committed 24,559 lines of code within the 188 days between the initial commit and the first release. As a naive, but reasonable approximation, we assumed this period reflects the minimal required effort to build a visualization tool. Thus, implementing separate visualization tools for the studied DBMSs would require around $188*5=940$ days. Using UPlan, we modified only around 800 lines of code of PEV2 to support all DBMSs. For PEV2, the average development speed is $24,559/188\approx 130$ lines of code per day, so the estimated effort for adapting UPlan is $800/130\approx 6$ days. Compared to implementing DBMS-specific visualization tools, which would require 940 days, UPlan would take only $194$ days ($188+6$) and would reduce the time effort by approximately 80%. Furthermore, the percentage of effort reduction would increase as the number of supported DBMSs grows. Note that this approximation ignores many factors, such as the complexity of query plans and the time to understand visualization tools.